# Deployability-Centric Infrastructure-as-Code Generation: An LLM-based Iterative Framework

### v1

@@ III-C Comparison with IaC-Eval Benchmark
Both DPIaC-Eval and IaC-Eval [9] cover approximately 60 AWS services, though each includes unique services not found in the other. DPIaC-Eval presents greater challenges than existing IaC benchmarks, with longer average line of code (155 v.s. 42) and more resources per template (7 v.s. 4). More importantly, DPIaC-Eval provides comprehensive coverage in four critical dimensions: syntax correctness, deployability, user intent matching, and security. In contrast, IaC-Eval supports only syntax and partially user intent matching. This multi-dimensional validation approach enables our benchmark to identify critical gaps in generated IaC template that would remain undetected in syntax-only evaluations. As we report in Section VI, the deployment validation reveals 42.7% of the syntactically correct templates fail during deployment, highlighting the importance of testing beyond mere syntax. Similarly, the security validation identifies only 9.0% of the deployable templates meet security compliance standards, exposing a critical shortcoming that is overlooked by IaC-Eval.
III-D Formative Study
Figure 2: Pass@1 scores across difficulty levels.
@@ III-D Formative Study
We first conduct a simple evaluation of mainstream LLMs on IaC template generation task by DPIaC-Eval benchmark. Fig. 2 presents the pass@1 accuracy results for six prominent LLMs across five difficulty levels. The results reveal that overall performance demonstrates significant room for improvement, with even the best-performing models achieving only sightly above 50% deployment success rates in difficulty level 1. Notably, all models exhibit a consistent inverse relationship between task complexity and performance, with accuracy declining substantially as difficulty increases from Level 1 to Level 5. The average pass@1 at difficulty level 5 is just 12.8%. This pattern indicates that current LLMs struggle with the complexity of real-world infrastructure provisioning scenarios. Thus, the preliminary results motivate us to propose a framework to fulfill this imperative research gap.
Figure 3: The overview of our framework. The number in Dataset Construction is the number of templates passed in from each action.${}^{1}$: a local variable in Python code.${}^{2}$o: indicates validation failure.${}^{3}$es: shows validation success.${}^{4}$ail too many times: represents exceeding the maximum iteration threshold.
IV Framework
LLMs have been widely adopted in various code and program generation tasks [32, 33, 34, 35, 36], achieving state-of-the-art performance. Current LLM approaches commonly treat IaC template generation as a static task [9, 16], failing to capture the iterative refinement process that characterizes real-world IaC development and consequently achieving poor deployment success. Our framework, IaCGen, addresses this limitation by implementing a deployment-driven approach that enables LLMs to learn from deployment failures and iteratively improve templates until they achieve successful deployment. As shown in Fig. 3, IaCGen consists of two modules, Iterative IaC Generation and IaC Deployment, integrated through a tiered feedback mechanism that mirrors real DevOps workflows. Due to the page limit, we focus on describing the high-level components and the complete set of prompts are available in our code repository.
IV-A Iterative IaC Generation
@@
Briefly, Iterative IaC Generation takes an IaC problem description, written in natural language, as input and produces IaC templates as output through a conversational feedback loop. The process works as follows and the numbers are indicated in Fig. 3: 1 A problem description prompt is retrieved from the dataset and fed to the LLM. 2 The LLM generates an initial IaC template based on the problem description. 3 The generated template and original prompt are stored in memory (implemented as a variable named conversation_history in code). 4 This template is passed to the IaC Deployment module for multi-stage validation. 5 If errors occur in IaC Deployment module, adaptive feedback is generated based on the specific failure point. We will elaborate on the adaptive feedback in Section IV-C. 6 This feedback is stored in the conversation history along with previous interactions. 7 The LLM is provided with both the feedback and all previous conversation history to generate an updated IaC template. Steps 2 to 7 are repeated iteratively until either the IaC template successfully passes all three validation stages or the LLM exceeds the predetermined maximum number of allowed attempts.
IV-B IaC Deployment
@@
This module takes the LLM-generated IaC template as input and validates its soundness through a three-stage deployment. Similar to manual deployment, this process typically consists of sequential validations, from format verification to schema checking to actual live deployment, with engineers iteratively refining their templates based on error messages [37, 38, 39]. If an error occurs at any stage, feedback is generated and returned to the Iterative IaC Generation module. If the template passes all three validation stages, the framework will end with the current IaC template as the final output (step 8).
@@
Format Verification. Most IaC solutions use declarative definition files to specify infrastructure configurations. The syntax for describing IaC usually depends on the requirements of the target platform, with common formats including YAML and JSON. The first validation stage verifies whether the LLM-generated IaC template conforms to the correct format specification. As AWS CloudFormation templates written in YAML, we employ yamllint [40] to validate basic structural correctness, including proper indentation, key-value pair formatting, and quotation usage.
@@
Schema Checking. After confirming proper formatting, the second validation stage uses a static analysis tool, cfn-linter [29], to verify if the code is correct in CloudFormation syntax. Although most existing LLM-based IaC frameworks stop after this stage, our approach extends validation to actual deployment testing, which addresses a critical gap in operational viability assessment.
@@
Live Deployment:. IaC shares a critical characteristic with traditional programming: code that passes compilation or syntax validation does not necessarily execute successfully at runtime. However, this challenge is particularly acute with IaC because of its unique properties. Most IaC languages are declarative rather than imperative, specifying the desired end state rather than procedural steps. This fundamental difference creates a validation gap that static analysis alone cannot bridge [41]. Cloud environments impose numerous constraints that syntax validators cannot fully capture: regional service availability variations, deprecated resource identifiers, and service interdependencies. For example, a syntactically perfect Terraform template referencing an outdated Amazon Machine Image ID will pass linting but fail during actual provisioning. Thus, we employ live deployment validation with the AWS SDK boto3 client, the official toolkit developed by AWS.
IV-C The Iterative Feedback
@@
Previous research have shown the effectiveness of feedback mechanisms in code generation tasks [42, 43, 44, 45]. IaC deployment inherently involves continuous integration and development with cycles of validation and refinement [4, 46]. By integrating a feedback loop mechanism into our framework, we simulate this real-world IaC development workflow while treating LLMs as “human developers”, evaluating not only an LLM’s initial generation capability, but also its error-resolution capacity. Our feedback design implements a progressive assistance approach that mirrors established DevOps practices for IaC development. This iterative feedback reflects how cloud engineers typically engage with increasingly specific guidance when resolving deployment errors.
@@
General Feedback. It provides high-level error indication by pointing out which validation stage the IaC template failed during the multi-stage validation process described above. For example, when a YAML syntax error is detected, the LLM might receive: “Based on the evaluation, the template contains YAML Syntax Errors.” as feedback. This mirrors the basic error alerts that developers encounter in CI/CD pipelines, requiring the LLM to diagnose and resolve issues with minimal guidance.
@@
Detailed Feedback. It offers more specific guidance by including complete error messages returned by the validation tools (yamllint, cfn-lint, and AWS SDK). For instance, instead of simply indicating a deployment error, the LLM receives the full error message: “An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [KeyName] must have values.” This level of feedback approximates the detailed diagnostic information that developers typically encounter when working with IaC testing tools [47] and tests the LLM’s ability to interpret technical error information. Considering the experimental efficiency and practicability, we constrain feedback loops to a maximum of six iterations per stage, with two attempts for general feedback and four attempts for detailed feedback. The asymmetrical allocation strategy is consistent with the empirical observations and experimental design in previous work [48].
V Experimental Settings
@@
To comprehensively evaluate the performance of IaCGen, we conduct a thorough multidimensional evaluation guided by the following research question:
@@
  • RQ2 What are the primary barriers preventing IaCGen from generating deployable IaC templates?
@@
  • RQ4 How trustworthy is the IaC templates generated by IaCGen?
TABLE II: Model performance with iterative feedback. A column with an upward arrow indicates its score is higher than the average.
Model passItr@1 passItr@5 passItr@10 passItr@15
@@
We select six representative leading LLMs, DeepSeek-S3, DeepSeek-R1, GPT-4o, GPT-o3-mini, Claude-3.5, and Claude-3.7. Those models cover both open-sourced and closed-sourced ecosystems, broadly demonstrating competitive performance in general code generation tasks [13, 49, 50, 51] and have significant industry adoption [52, 53]. All evaluations are conducted through the APIs from official providers ensuring reproducibility and standardization. All models are evaluated with consistent temperature settings of 0 to maximize deterministic outputs and configured with the 8,000 maximum output token limit to accommodate complex IaC templates. We run those models on the benchmark introduced in Section III.
To ensure unbiased deployment evaluation, we create an AWS sandbox account with default service quota and no pre-existing resources. Each validation cycle follows a complete life cycle, i.e., allocating resources, validating their operational state, and then structurally removing all resources to set the environment back to its initial state. This approach ensures that each test runs in a clean environment, preventing “cross-contamination” between evaluation runs.
Evaluation Metrics. The unbiased version of pass@k [13] is a standard metric used by current benchmarks [13, 54, 14], which indicates the probability that at least one of the $k$ generated samples yields a correct solution. However, it does not effectively capture the performance of iterative feedback. To address this limitation, we introduce passItr@n, where $n$ represents the maximum number of iterations allowed. passItr@n indicates the probability that the IaC template is correctly generated before or at iteration $n$.
VI Evaluation
VI-A RQ1: How effective is IaCGen in IaC template generation?
@@
Table II presents the performance of our framework on the DPIaC-Eval benchmark as the iteration increases. The initial low success rates (passItr@1) align with the previous finding [9], with an average score of approximately 24.7% on the first attempt of all models. This demonstrates the inherent challenge of LLM-based IaC template generation compared to general-purpose code generation tasks, where the average score of the models is 95.6% in HumanEval benchmark [13].
@@
Furthermore, all models achieve better performance with more iterations. With sufficient iterations, all models eventually achieve a score higher than 50%, and average of 75.2% pass@Itr15 accuracy, which is a near 200% performance improvement as compared to pass@Itr@1. This shows IaCGen exhibits capability on effectively utilizing feedback to correct errors in IaC templates. Claude models show impressive performance in the generation of IaC templates, achieving leading results of over 90% passItr@15 accuracy among the models assessed.
@@
Iterative Feedback. Fig. 4 shows the improvements with the iterative feedback mechanism for all models. We observe that all models moderately benefit from general feedback (4.6% to 17.6%) and largely improve from detailed feedback (28.1% to 51.0%), highlighting that models can effectively learn from the iterative feedback. Claude models demonstrate remarkable performance, achieving a high passItr@15 score (95.5% for Claude-3.5 and 92.8% for Claude-3.7) using general and detailed feedback. In contrast, GPT-4o shows substantially lower performance (55.2%).
Figure 4: Distribution of feedback levels among successfully generate records.
@@
Effectiveness of Conversation History. Unlike previous research [9, 48, 16, 55] that isolates each feedback iteration, our approach preserves the entire conversation history throughout the iterative feedback. To evaluate the effectiveness of our iterative feedback mechanism that provides LLMs with the entire conversation history, we conduct an ablation study that compares it to the traditional approach, which only feeds the system prompt, original prompt, error message, and incorrect IaC template to the LLM. We select Claude-3.5, the best performing model, for this study. As illustrated in Fig. 5, both approaches demonstrate comparable performance during the first two iterations. However, our approach establishes a consistent advantage starting with the third iteration. Our method achieves a deployable IaC template in an average of 4.55 iterations, while the traditional approach required an average of 5.41 iterations, which is an 15.9% reduction in required iterations. Moreover, without accessing the entire conversation history, we find that LLMs frequently reintroduce previously corrected errors when addressing new issues. This “Error Recurrence” phenomenon suggests that maintaining the conversation history provides a crucial context that enables the LLMs to learn from past corrections and avoid repeating similar mistakes across iterations.
Figure 5: PassItr@n score comparison for IaCGen and normal LLMs without iterative feedback on Claude-3.5.
Answer to RQ1: IaCGen demonstrates high effectiveness in IaC template generation. Through iterative feedback, it transforms the initial poor success rates ($\approx$24.7%) into high deployment capabilities (average 75.2%) for all evaluated models, and Claude models achieve over 90% passItr@15.
VI-B RQ2: What are the primary barriers preventing IaCGen from generating deployable IaC templates?
Fig. 6 shows when LLM encounters an error during IaC generation in all models evaluated. Syntax issues, which occur during the schema checking stage, constitute the primary challenge for nearly all models, followed by deployment errors that only manifest during live deployment testing. This finding validates our methodological focus on the validation of deployability beyond mere syntactic correctness, highlighting the gap between code that appears valid and code that functions correctly in production environments.
OpenAI’s models exhibit distinct error patterns compared to other models in our evaluation. While GPT-4o’s primary challenge is deployment errors (suggesting strong syntax capabilities but weaker understanding of operational constraints), GPT-o3-mini experiences a disproportionate rate of formatting errors ($>$30% compared to approximately 10% for other models). This discrepancy points to fundamental differences in how various model architectures approach structured template generation, potentially reflecting differences in training data composition or model architecture.
We identify five primary error categories with notable frequency and impact on model performance. Table III illustrates the error messages and the distribution among models.
TABLE III: Five most common errors across three validation stages. xx1: placeholder for line number and value name. For Count, from left to right are: GPT-4o, o3-mini, Claude-3.5, Claude-3.7, DeepSeek-S3, DeepSeek-R1.
Error Message Percentage Failed Stage Count
1 An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [xx1] must have values 14.69% deployment [Uncaptioned image]
2 Additional properties are not allowed (‘xx’ was unexpected) 6.44% syntax [Uncaptioned image]
3 ‘Fn::Sub’ isn’t needed because there are no variables 3.09% syntax [Uncaptioned image]
4 Line xx: too many spaces inside brackets 2.92% format [Uncaptioned image]
5 Parameter validation failed: parameter value xx for parameter name xx does not exist 1.19% deployment [Uncaptioned image]
@@ VI-B RQ2: What are the primary barriers preventing IaCGen from generating deployable IaC templates?
Figure 6: Error stages distributions across models.

### v2

@@ 3.3. Comparison with IaC-Eval Benchmark
Both DPIaC-Eval and IaC-Eval (kon2024iac) cover around 60 AWS services, though each includes unique services not found in the other. DPIaC-Eval presents greater challenges than existing IaC benchmarks, with longer average line of code (155 vs. 42) and more resources per template (7 vs. 4). More importantly, DPIaC-Eval provides comprehensive coverage in four critical dimensions: syntax correctness, deployability, user intent matching, and security. In contrast, IaC-Eval supports only syntax validation and user intent matching. This multi-dimensional validation approach enables our benchmark to identify critical gaps in generated IaC template that remain undetected in syntax-only evaluations. As we report in Section 6, the deployment validation reveals 42.7% of the syntactically correct templates fail during deployment, highlighting the importance of testing beyond mere syntax. Similarly, the security validation identifies only 9.0% of the deployable templates meet security compliance standards, exposing a critical shortcoming that is overlooked by IaC-Eval.
3.4. Formative Study
Figure 4. Pass@1 scores across difficulty levels.
@@ 3.4. Formative Study
We first conduct a simple evaluation of mainstream LLMs on IaC template generation task using DPIaC-Eval benchmark. Fig. 4 presents the pass@1 accuracy results for six prominent LLMs across five difficulty levels. The results reveal that overall performance has significant room for improvement, with the best-performing models achieving only sightly above 50% deployment success rates in level 1 difficulty. Notably, all models exhibit a consistent inverse relationship between task complexity and performance, with score declining as difficulty increases from Level 1 to Level 5. The average pass@1 at difficulty level 5 is just 12.8%. This pattern indicates that current LLMs struggle with the complexity of real-world infrastructure provisioning scenarios. Thus, the preliminary results motivate us to propose a framework to fulfill this imperative research gap.
4. Framework
LLMs have achieved remarkable success in code generation tasks (lachaux2020unsupervised; roziere2021leveraging; si2024solution; pan2024large), demonstrating state-of-the-art performance across programming languages. However, their application to IaC generation remains limited and problematic. Current approaches (kon2024iac; ragothaman2024optimizing; pujar2023automated) treat IaC templates as conventional programs, focusing primarily on syntactic correctness while neglecting their continuous integration and deployment (CI/CD) nature (srivatsa2024survey; iac-ci-cd). Consequently, generated templates often fail during actual deployment despite being syntactically valid and resulting in significantly lower accuracy rates compared to general programming tasks (GPT 4 achieves 19% in IaC-Eval and 83% in HumanEval). Our framework, IaCGen, addresses this limitation through a deployment-driven approach that enables LLMs to learn from deployment failures and continuously refine templates until achieving successful deployment. As shown in Fig. 5, IaCGen consists of two modules, Iterative IaC Generation and IaC Deployment, integrated through a tiered feedback mechanism that mirrors real DevOps workflows. The complete prompts and other reproducible details mentioned in this section are available in our code repository.
4.1. Iterative IaC Generation
@@
Briefly, Iterative IaC Generation takes an IaC template description, written in natural language, as input and produces IaC template as output through a conversational feedback loop. The process works as follows and the numbers are indicated in Fig. 5: 1 A natural language prompt about the infrastructure requirement is fed to the LLM. 2 The LLM generates an initial IaC template based on the requirement. 3 The generated template and original prompt are stored in memory (implemented as a variable named conversation_history in code). 4 This template is passed to the IaC Deployment module for multi-stage validation. 5 If errors occur in IaC Deployment module, adaptive feedback is generated based on the specific failure point. We will elaborate on the adaptive feedback in Section 4.3. 6 This feedback is stored in the conversation history along with previous interactions. 7 The LLM is provided with both the feedback and all previous conversation history to generate an updated IaC template. To ensure output quality and consistency, in step 1, we apply a system prompt containing role assignment and task description, along with a lightweight chain-of-thought strategy to guide the generation process. Steps 2 to 7 are repeated iteratively until either the IaC template successfully passes all three validation stages or the LLM exceeds the predetermined maximum number of allowed attempts.
[…]
4.2. IaC Deployment
@@
This module takes the LLM-generated IaC template as input and validates its soundness through a three-stage deployment. Similarly to real-world testing, this process consists of sequential validations, from format verification to syntax checking to live deployment, with LLM iteratively refining the template based on error messages (iac-best-practices; vessey1985expertise; debugging-web). If an error occurs at any stage, feedback is generated and returned to the Iterative IaC Generation module. If the template passes all three validation stages, the framework will end with the current IaC template as the final output (step 8).
@@
Format Verification. Most IaC solutions use declarative definition files to specify infrastructure configurations. The syntax for describing IaC usually depends on the requirements of the target platform, with common formats including YAML and JSON. The first validation stage verifies whether the LLM-generated IaC template conforms to the correct format specification. As AWS CloudFormation templates written in YAML, we employ yamllint (yaml-linter) to validate basic structural correctness, including proper indentation, key-value pair formatting, and quotation usage.
@@
Syntax Checking. After confirming proper formatting, the second validation stage uses a static analysis tool, cfn-linter (cfn-linter), to verify if the code is correct in CloudFormation syntax. Although most existing LLM-based IaC frameworks stop after this stage, our approach extends validation to actual deployment testing, which addresses a critical gap in operational viability assessment.
@@
Live Deployment. IaC shares a critical characteristic with traditional programming: code that passes compilation or syntax validation does not necessarily execute successfully at runtime. However, this challenge is particularly acute with IaC because of its unique properties. Most IaC languages are declarative rather than imperative, specifying the desired end state rather than procedural steps. This fundamental difference creates a validation gap that static analysis alone cannot bridge (guerriero2019adoption). Cloud environments impose numerous constraints that syntax validators cannot fully capture: regional service availability variations, deprecated resource identifiers, and service interdependencies. For example, a syntactically perfect Terraform template referencing an outdated Amazon Machine Image ID will pass linting but fail during actual provisioning. Thus, we employ live deployment validation with the AWS SDK boto3 client, the official toolkit developed by AWS.
4.3. The Iterative Feedback
@@
Previous research have shown the effectiveness of feedback mechanisms in code generation tasks (chen2023teaching; yang2022generating; huang2022large; zhou2024bridging). IaC deployment inherently involves continuous integration and development with cycles of validation and refinement (sokolowski2022infrastructure; aws-best-practices). By integrating a feedback loop mechanism into our framework, we simulate this real-world IaC development workflow while treating LLMs as “human developers”, evaluating not only an LLM’s initial generation capability, but also its error-resolution capacity. Our feedback design implements a progressive assistance approach that mirrors established DevOps practices for IaC development. This iterative feedback reflects how cloud engineers typically engage with increasingly specific guidance when resolving deployment errors.
@@
General Feedback. It provides high-level error indication by pointing out which validation stage failed during the multi-stage validation process described above. Feedback follows a standardized structure:“Based on the evaluation, the template contains [type of error] Errors.” The types of error are YAML Syntax, CloudFormation Template Syntax, and Deployment. For example, when errors occur during the “Formate Verification” stage, the LLM receives: “Based on the evaluation, the template contains YAML Syntax Errors.”. This mirrors the basic error alerts that developers encounter in CI/CD pipelines, requiring the LLM to diagnose and resolve issues with minimal guidance.
@@
Detailed Feedback. It offers more specific guidance by including complete error messages returned by the validation tools (yamllint, cfn-lint, and AWS SDK). For instance, instead of simply indicating a deployment error, the LLM receives the full error message: “An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [KeyName] must have values.” This level of feedback approximates the detailed diagnostic information that developers typically encounter when working with IaC testing tools (hasan2020testing) and tests the LLM’s ability to interpret technical error information. Considering the experimental efficiency and practicability, we constrain feedback loops to a maximum of six iterations per stage, with two attempts for general feedback and four attempts for detailed feedback. The asymmetrical allocation strategy is consistent with the empirical observations and experimental design in previous work (palavalli2024using).
5. Experimental Settings
@@
To comprehensively evaluate the performance of IaCGen, we conduct a thorough multidimensional evaluation guided by the following research questions:
@@
  • RQ2 What are the primary barriers of generating deployable IaC templates?
@@
  • RQ4 How trustworthy are the IaC templates generated by IaCGen?
We select six representative leading LLMs, GPT-4o, GPT-o3-mini, Claude-3.5, Claude-3.7, DeepSeek-R1, and DeepSeek-V3. Those models cover both open-sourced and closed-sourced ecosystems, broadly demonstrating competitive performance in general code generation tasks (chen2021evaluating; sobo2025evaluating; islam2025codesim; guo2024deepseek) and have significant industry adoption (pan2024textit; deepseek-share). All evaluations are conducted through the APIs from official providers ensuring reproducibility and standardization. All models are evaluated with consistent temperature settings of 0 to maximize deterministic outputs and configured with the 8,000 maximum output token limit to accommodate complex IaC templates. We run those models on the benchmark introduced in Section 3.
To ensure unbiased deployment evaluation, we create an isolated AWS account with default service quota and no pre-existing resources to mimic real production environment. Each validation cycle follows: allocating resources, validating their operational state, and then structurally removing all resources to set the environment back to its initial state. This approach ensures that each test runs in a clean environment, preventing “cross-contamination” between evaluation runs.
Evaluation Metrics. The unbiased version of pass@k (chen2021evaluating) is a standard metric used by current benchmarks (chen2021evaluating; austin2021program; liu2023your), which indicates the probability that at least one of the $k$ generated samples yields a correct solution. However, it does not effectively capture the performance of iterative feedback. To address this limitation, we introduce passItr@n, where $n$ represents the maximum number of iterations allowed. PassItr@n indicates the success rate that the IaC template is correctly generated (able to be deployed successfully) before or at iteration $n$.
Experiment costs. The cost of the study consists of two parts, one for the LLM API Token and one for the Deployment validations on AWS. The total cost (USD) is \$230.75 and \$35.21, respectively. On average, the token cost of the most expensive model (Claude-3.7-Sonnet) is \$0.42, but the cheapest model (DeepSeek-V3) is \$0.04 only, and the deployment cost is \$0.04, for each deployable template. These modest costs demonstrate that IaCGen provides a cost-effective approach for practical implementation with the lowest of \$0.08 to generate one deployable template. The AWS CloudFormation requires no fee for the deployment test. However, the deployment of some resources require fees once it is deployed such as the Relational Database Service.
6. Evaluation
Table 2. Model performance with iterative feedback. A column with an upward arrow indicates its score is higher than the average. pItr stands for passItr.
Model pItr@1 pItr@5 pItr@10 pItr@15
@@ 6. Evaluation
Figure 6. Distribution of maximum feedback level required for successfully generated templates. ds stands for deepseek.
6.1. RQ1: How effective is IaCGen in IaC template generation?
@@
Table 2 presents the performance of our framework on the DPIaC-Eval benchmark as the iteration increases. The initial low success rates (passItr@1) align with the previous finding (kon2024iac), with an average score of approximately 24.7% on the first attempt of all models. This demonstrates the inherent challenge of LLM-based IaC template generation compared to general-purpose code generation tasks, where the average score of the models is 95.6% in HumanEval benchmark (chen2021evaluating).
@@
Furthermore, all models achieve better performance with more iterations. With sufficient iterations, all models eventually achieve a score higher than 50%, and average of 75.2% pass@Itr15 accuracy, which is a near 200% performance improvement as compared to passItr@1. This shows IaCGen exhibits capability on effectively utilizing feedback to correct errors in IaC templates. Claude models show impressive performance in the generation of IaC templates, achieving leading results of over 90% passItr@15 accuracy among the models assessed.
@@
Iterative Feedback. Fig. 6 shows the improvements with the iterative feedback mechanism for all models. We observe that all models moderately benefit from general feedback (4.6% to 17.6%) and largely improve from detailed feedback (28.1% to 51.0%), highlighting that models can effectively learn from the iterative feedback. Claude models demonstrate remarkable performance, achieving a high passItr@15 score (95.5% for Claude-3.5 and 92.8% for Claude-3.7) In contrast, GPT-4o shows substantially lower performance (55.2%).
@@
Effectiveness of Conversation History. Unlike previous research (kon2024iac; palavalli2024using; ragothaman2024optimizing; madaan2023self) which isolates each feedback iteration, our approach preserves the entire conversation history for each iteration. To evaluate the effectiveness of our iterative feedback mechanism that provides LLMs with the entire conversation history, we conduct an ablation study that compares it to the baseline approach, which only feeds the system prompt, original prompt, error message, and incorrect IaC template to the LLM. We select Claude-3.5, the best performing model, for this study. As illustrated in Fig. 8, both approaches demonstrate comparable performance during the first two iterations. However, our approach establishes a consistent advantage starting with the third iteration. Our method achieves a deployable IaC template in an average of 4.55 iterations, while the baseline approach required an average of 5.41 iterations, which is an 15.9% reduction in required iterations.
Without accessing the entire conversation history, we find that LLMs frequently reintroduce previously corrected errors when addressing new issues. This “Error Recurrence” phenomenon suggests that maintaining the conversation history provides a crucial context for LLMs to learn from past corrections and avoid repeating similar mistakes across iterations. This improvement might stem from IaC’s inherent complexity, which deployment errors often result from intricate resource interdependencies rather than isolated issues, yet error messages typically identify only the immediate failure point. By providing complete conversation history, LLMs can understand the cascading effects of their previous modifications and make more informed corrections. This contextual awareness enables systematic problem-solving rather than random attempts, as evidenced by the reduced error recurrence and faster convergence to deployable templates.
Figure 7. PassItr@n scores of IaCGen with and without conversation history on Claude-3.5.
@@
Figure 8. Distribution of error stages (i.e., where the generated template fails) across models.
Generalizability of IaCGen. IaCGen’s modular design and the cross-language capabilities of LLMs enables seamless adaptation to different IaC technologies. Practitioners can adapt IaCGen to alternative IaC languages and cloud service providers by substituting the corresponding linters (for format and syntax validation) and deployment components (for deployment verification). To evaluate IaCGen’s generalizability quantitatively, we conduct experiments on the IaC-Eval benchmark (kon2024iac), which contains Terraform templates, with updates to the linter. Due to fundamental architectural differences between CloudFormation and Terraform (discussed in Section 2.2), we focus on syntax validation following the established methodology of IaC-Eval, using Claude-3.5 based on its superior performance in our primary experiments. IaCGen demonstrates strong generalizability to Terraform, achieving 79.7% passItr@1 and 100% passItr@7 accuracy with an average of 1.58 iterations. For comparison, IaCGen requires 2.53 iterations on average to achieve 100% passItr@5 accuracy on DPIaC-Eval’s CloudFormation templates. This difference reflects the higher complexity of DPIaC-Eval templates, evidenced by their higher average lines of code and resource complexity.
Answer to RQ1: IaCGen demonstrates high effectiveness in IaC template generation. Through iterative feedback, it transforms the initial poor success rates ($\approx$24.7%) into high deployment capabilities ($\approx$75.2%) for all evaluated models, and Claude models achieve over 90% passItr@15. IaCGen also shows its strong generalizability to Terraform with $\approx$25% performance improvement. Maintaining complete conversation history can effectively mitigate the “Error Recurrence” phenomenon.
6.2. RQ2: What are the primary barriers of generating deployable IaC templates?
Fig. 8 shows the distribution of error stage where LLMs encounter the first deployability error during IaC generation. Syntax issues, which occur during the syntax checking stage, constitute the primary challenge for nearly all models, followed by deployment errors that only manifest during live deployment testing. This finding validates our methodological focus on the validation of deployability beyond mere syntactic correctness, highlighting the gap between code that appears valid and code that functions correctly in production environments.
OpenAI’s models exhibit distinct error patterns compared to other models in our evaluation. While GPT-4o’s primary challenge is deployment errors (suggesting strong syntax capabilities but weaker understanding of operational constraints), GPT-o3-mini experiences a disproportionate rate of formatting errors (>>30% compared to approximately 10% for other models). This discrepancy points to fundamental differences in how various model architectures approach structured template generation, potentially reflecting differences in training data composition or model architecture.
We identify five primary error categories with notable frequency and impact on model performance. Table 3 illustrates the error messages and the distribution among models.
Table 3. Five most common errors across three validation stages. xx1: placeholder for line number and value name. For Count, from left to right are: GPT-4o, o3-mini, Claude-3.5, Claude-3.7, DeepSeek-R1, DeepSeek-V3.
# Error Type Error Message Percentage Failed Stage Count
1 Missing Value An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [xx1] must have values 14.69% Deployment [Uncaptioned image]
2 Self-defined Property Additional properties are not allowed (‘xx’ was unexpected) 6.44% Syntax [Uncaptioned image]
3 Null Substitution ‘Fn::Sub’ isn’t needed because there are no variables 3.09% Syntax [Uncaptioned image]
4 Unnecessary Whitespace Line xx: too many spaces inside brackets 2.92% Format [Uncaptioned image]
5 Arbitrary Default Value Parameter validation failed: parameter value xx for parameter name xx does not exist 1.19% Deployment [Uncaptioned image]